# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 2.2 Conversational Engagement
Conversational engagement tasks in multi-turn dialogue have been catalyzed by several pioneering studies that established how to define and measure an LLM’s capacity to sustain interactions over multiple turns. One early effort is the ABC-Eval framework by Finch et al. (2023) [19], which introduced a dimensional human evaluation scheme for open-domain chat. ABC-Eval defines 16 fine-grained conversational behavior categories (spanning aspects like factual accuracy, consistency, relevance, and empathy) and uses these as binary turn-level labels to quantify dialogue quality. Although this benchmark relies on labor-intensive human evaluations, its initial findings underscore the inherent challenges in assessing iterative interactions, thereby motivating the development of more scalable and nuanced evaluation frameworks.
Shortly afterward, Duan et al. (2023) [25] introduced the BotChat evaluation paradigm, which reduces reliance on costly human judges by leveraging LLMs for both conversation generation and evaluation. In BotChat, models are prompted to extend real-world dialogue seeds (ChatSEED prompts) into extended multi-turn conversations, subsequently evaluated by top-tier LLMs (e.g., GPT-4) serving as automated judges. Notably, GPT-4-generated dialogues were found to be nearly indistinguishable from human conversations, successfully fooling discriminator models, whereas other contemporary LLMs exhibited shortcomings in instruction adherence and conciseness, underscoring specific challenges in maintaining human-like coherence across multi-turn dialogues.
Most recently, Sirdeshmukh et al. (2025) [72] introduced MultiChallenge, a benchmark designed to rigorously test conversational persistence and context management in frontier LLMs. MultiChallenge comprises four realistic scenarios—long-term instruction retention, implicit information recall, iterative revision, and consistent responses without sycophancy—requiring simultaneous instruction-following, context tracking, and reasoning.
@@ 2.2.1 Conversational Engagements in Roleplay
Early persona-grounded dialogue systems aimed at maintaining consistent character personas over multi-turn conversations using architectures like memory networks and transformers. Significant contributions included Li et al. (2016) [73] introducing persona embeddings, Kottur et al. [74] using personalized memory networks, and notably, Zhang et al. (2018) [75]’s PersonaChat dataset and memory-based models that set foundational benchmarks for persona consistency. These initial works primarily trained models from scratch, facing challenges in sustained persona adherence across interactions. The recent survey [76] extensively covers role-playing before and after the advent of LLMs. Within the scope of this survey, we focus specifically on role-playing interactions under multi-turn LLM settings.
@@
Early LLMs demonstrated an ability to impersonate user-defined personas through prompting alone. Users could supply descriptions like “You are a wise old wizard…” in a system or context prompt, and models like GPT-3 would attempt to respond “in character.” PersonaLLM [77] introduced a benchmark for personalization and highlighted that simply prefixing instructions with high-level persona descriptions yields limited diversity; instead, it proposed simulating nuanced user preferences via prompt-based reward models. It showed that prompting can move beyond trivial traits to tailor outputs to idiosyncratic user needs. Similarly, CharacterChat [78] used role-playing prompts with behavior presets and dynamic memory to maintain a character’s persona over long conversations. It constructed an MBTI-based persona bank and prompted ChatGPT to produce dialogues between a “seeker” and a compatible “supporter,” injecting preset behavioral tendencies and retrieving context-specific memory each turn to keep interactions coherent.
Beyond persona style, prompting has been used to improve reasoning. Role-play prompting [79] showed that instructing an LLM to “pretend to be” a domain expert can implicitly trigger step-by-step reasoning. In zero-shot settings across 12 tasks, prompting models with a role (e.g. “You are an excellent math teacher…”) led to significantly higher accuracy than a vanilla prompt. The method uses a two-stage prompting framework: first having the model generate an “immersive” backstory or persona acknowledgement, then using that along with the query. These studies collectively demonstrate prompting-based role-play as a powerful lever: it can induce consistent persona adherence (persona and behavior presets) and even boost cognitive performance (reasoning via expert roles) without any parameter updates.
@@
To achieve more robust in-character behavior, researchers introduced instruction tuning and fine-tuning with role-playing data. PIPPA [80] released a partially synthetic corpus of over 1 million role-play messages, crowdsourced from an online community of role-play enthusiasts. By fine-tuning on PIPPA’s diverse persona-conditioned conversations, small LLMs dramatically improved at staying in character, underscoring that sheer volume and diversity of persona-rich dialogues can teach consistent role-playing. UltraChat [81] constructed 1.5M multi-turn dialogues via self-chat with GPT-4, covering broad topics and user types. Fine-tuning LLaMA on this yielded UltraLLaMA, which surpassed previous open models in general conversation quality (including user engagement and coherence). Other data efforts target specific role-play domains: PRODIGy [82] built a dialogue dataset from movie scripts aligned with detailed character profiles (biographies, personality traits). Fine-tuning models on these profile-grounded movie dialogues significantly improved consistency when emulating those characters – e.g. including a character’s backstory and speaking style led to higher human preference for in-character responses.
@@
A parallel direction creates models specialized for particular characters or customizable personas. ChatHaruhi [83] focused on anime characters, compiling 54k dialogues for 32 characters by combining original script lines with simulated conversations. The model fine-tuned on this data, augmented with a “memory” of past events for each character, was able to “revive” characters like Haruhi Suzumiya, accurately quoting lore and personality in new interactions.In another example, CharacterGLM [84] built on the Chinese GLM model to allow explicit profile injection for any character: it fine-tuned ChatGLM variants on a corpus of dialogues with richly annotated character profiles (covering identity, style, relationships) and achieved state-of-the-art human-likeness and consistency for customized personas. RoleCraft-GLM [85] further extended this concept by crafting original non-celebrity personas with emotional depth and fine-tuning a model on dialogues involving those characters. This yielded more nuanced emotional consistency, validating that meticulous character development during fine-tuning yields agents that are engaging and lifelike in their persona.
@@
Recent work also explores how to fine-tune effectively for role consistency. Ditto [86] introduced a self-alignment pipeline where the model generates its own role-play dialogues for 4,000 distinct characters and then trains on them. By leveraging the model’s internal knowledge to create training data (with feedback to ensure each character stays distinct), Ditto achieved strong persona fidelity across a wide range of roles, outperforming other fine-tuned baselines on a Role-play benchmark. This highlights a trend of using LLMs themselves to amplify training: both UltraChat and Ditto use models to generate synthetic data, but Ditto’s focus is specifically on persona diversity and consistency (it treats role-play generation as a reading-comprehension task to avoid style collapse, then fine-tunes on the result). Another notable work, CharacterLLM [87], demonstrated a pipeline where for each target character (especially historical figures), one uses Wiki bios and documents to prompt an LLM to produce dialogues involving that character. Fine-tuning on this synthetic dialogue allowed a single agent to robustly play many famous roles, effectively transferring factual knowledge into conversational skill. In summary, supervised fine-tuning for role-play has evolved from harvesting large-scale data to increasingly clever data generation and specialization techniques.
@@
Even with instruction tuning, a given model has limits in the number of distinct characters or styles it can perfectly emulate. Thus, a key theme is personalization: adapting an LLM to a new persona with minimal data or effort. Recent research has explored parameter-efficient tuning modules that allow rapid persona swapping without retraining the entire model. PersonaPKT [88] proposed representing each persona as a continuous embedding vector that can be learned from a small set of that user’s dialogues. By keeping the pre-trained model fixed and only training a tiny persona-specific vector (less than 0.1% of parameters), PersonaPKT efficiently imbues the model with that user’s speaking style and preferences. In a related vein, PPlug [89] introduced a plug-and-play user encoder that on-the-fly computes a user embedding from their conversation history . Instead of a static learned vector, it employs a small model to read a user’s past messages and output a “personal embedding” summarizing their quirks and facts. This embedding is then prepended to the LLM’s input to personalize the response. Such design lets the LLM be dynamically personalized each turn based on context, and experiments showed significant gains in personalization metrics across tasks.
As LLMs are used to power multiple characters simultaneously, it becomes important to switch personas quickly and even maintain many personas at once. Neeko [90] addresses this with a dynamic LoRA (Low-Rank Adapter) framework for multi-character role-play. It pre-trains a separate LoRA module for each character and uses a gating network to activate the appropriate one based on the dialogue context. This incremental ability means an unlimited number of personas can gradually accrue, each encapsulated in a plug-in adapter. An agent can seamlessly swap roles by toggling adapters, which showed superior consistency when one model needed to portray many characters in a group chat. Another challenge is preserving persona over multiple dialogue rounds. Standard fine-tuning often splits dialogues into independent turns, which can break character memory. MIDI-Tuning [91] proposed to explicitly model the user and system roles with separate adapters and a round-level state. In their framework, an LLM-based agent is trained by alternating between a “user adapter” (processing user utterances) and an “agent adapter” (generating responses), carrying hidden state forward through turns.
@@
While supervised learning can teach a model a persona, it doesn’t explicitly punish lapses. Reinforcement learning (RL) provides a way to directly optimize consistency and other long-horizon behaviors. Shea et al. (2023) [92] demonstrated this by applying offline RL to a dialogue model for persona consistency. They took an existing high-quality chatbot and defined a reward that penalizes persona breaks (e.g. contradicting provided profile or previous statements) and rewards in-character responses. Rather than interact with humans live, they performed RL on a static dataset of conversations – adjusting the model’s policy to maximize the persona consistency reward while leveraging off-policy data. The result was a chatbot that, in human evaluation, more reliably adhered to its given persona description and avoided contradictions compared to its purely supervised counterpart. This work bridged the gap between static fine-tuning and RLHF: it shows one can inexpensively refine a model to be more in-character by offline RL on existing dialogues, getting some benefits of RL (direct control of behavior) without an expensive online loop.
Another aspect of multi-turn role-play that benefits from RL-like thinking is maintaining long-term coherence. COMEDY [93] approached this via a compressive memory mechanism that can be seen as the model “reinforcing” important memory content over a conversation. Instead of a traditional retrieval pipeline, COMEDY has the LLM periodically summarize and compress the dialogue history (including user persona hints and past events) into a concise memo, which is fed back into itself for future responses. This one-model architecture learns through supervised fine-tuning to generate useful summaries (e.g. remembering the user’s preferences or the agent’s own backstory) and to consult them when answering. While not an RL in algorithm, COMEDY’s design implicitly optimizes a long-term reward: the compressed memory serves to avoid contradictions and boring repetition, much like an RL agent maximizing a reward for user engagement would learn to recall relevant facts.
@@
One line of work focuses on general personalization. The LaMP benchmark [94] evaluates whether LLMs can adapt to user-specific profiles across a range of tasks—such as rewriting text in a personalized tone or classifying based on individual preferences. While not limited to dialogue, LaMP highlights the broader need for systems that understand and leverage identity cues, and confirms that retrieval-based methods are especially effective for on-the-fly personalization.
@@
A more targeted category of benchmarks examines character-specific fidelity. CharacterEval [95] is a Chinese benchmark featuring 77 characters from novels, each with multi-turn dialogues and detailed profiles. It defines 13 metrics across four dimensions – including character consistency, behavior realism, and conversational quality – and provides human and model-based evaluation for each. CharacterEval revealed that even GPT-4, when role-playing in Chinese, could be outperformed by fine-tuned local models on consistency, indicating that specialized training made a measurable difference. On the English side, RoleEval [96] poses factual and commonsense questions about 300 well-known characters, assessing whether models accurately retain and apply character-specific knowledge. Results show that global models like GPT-4 excel with internationally known figures, while locally fine-tuned models do better with culturally specific roles. Meanwhile, TimeChara [97] explores a new dimension—temporal consistency—by checking if a model role-playing a character at a given point in a story inadvertently leaks future events. Even the strongest models often violate timeline boundaries, suggesting the need for narrative-aware mechanisms to constrain temporal knowledge. SimulBench [98] is a benchmark assessing LLM performance in interactive simulation scenarios – imaginative, role-playing and tool-use tasks that unfold over multiple turns. The benchmark includes tasks like acting as a Linux terminal, playing text-based games, and complex, long-horizon simulations requiring dynamic interaction with a user.
@@
Moving from factual to psychological evaluation, InCharacter [99] assesses if a role-playing agent truly internalizes a character’s personality. It uses an interview-style personality test: the agent (in role) is asked a battery of questions akin to a psychological survey, and its answers are compared to the expected personality profile of the character. Complementing this, RoleInteract [100] evaluates social interaction skills at two levels: one-on-one conversation quality (empathy, politeness, etc.) and group dynamics (how well an agent plays its role in a multi-agent conversation). RoleInteract includes 500 characters and diverse scenarios (e.g., an office meeting with several personas). One finding was that some agents that excel in bilateral chat struggled in group settings – sometimes a normally consistent character would conform or get sidetracked when other AI characters were present, indicating influence and social pressure effects.
@@
A different angle on evaluation is to test how well models understand a character from source material. Yuan et al. (2024) [101] argue that a truly aligned role-play model should be able to read a narrative and produce a coherent character profile. They created the CROSS dataset of expert-written character profiles (covering attributes, relationships, events, and personality) for characters in novels. Models are evaluated on how well they can generate similar profiles after “reading” the novel (or being given chapters as input). In addition to intrinsic metrics (overlap with the expert profile), they evaluate extrinsically via a motivation recognition task: given a scenario from the story, does the model’s profile help it answer why the character acted a certain way. Such evaluation drives home that role-play is not only about output style, but also about the model’s internal model of the character.

### v5

@@ 2.2 Conversational Engagement
Conversational engagement tasks in multi-turn dialogue have been catalyzed by several pioneering studies that established how to define and measure an LLM’s capacity to sustain interactions over multiple turns. One early effort is the ABC-Eval framework by Finch et al. (2023), which introduced a dimensional human evaluation scheme for open-domain chat. ABC-Eval defines 16 fine-grained conversational behavior categories (spanning aspects like factual accuracy, consistency, relevance, and empathy) and uses these as binary turn-level labels to quantify dialogue quality. Although this benchmark relies on labor-intensive human evaluations, its initial findings underscore the inherent challenges in assessing iterative interactions, thereby motivating the development of more scalable and nuanced evaluation frameworks.
Shortly afterward, Duan et al. (2024) introduce the BotChat evaluation paradigm, which reduces reliance on costly human judges by leveraging LLMs for both conversation generation and evaluation. In BotChat, models are prompted to extend real-world dialogue seeds (ChatSEED prompts) into extended multi-turn conversations, subsequently evaluated by top-tier LLMs (e.g., GPT-4) serving as automated judges. Notably, GPT-4-generated dialogues were found to be nearly indistinguishable from human conversations, successfully fooling discriminator models, whereas other contemporary LLMs exhibited shortcomings in instruction adherence and conciseness, underscoring specific challenges in maintaining human-like coherence across multi-turn dialogues.
Most recently, Deshpande et al. (2025) introduce MultiChallenge, a benchmark designed to rigorously evaluate conversational persistence and context management in frontier LLMs. MultiChallenge encompasses four realistic scenarios—long-term instruction retention, implicit information recall, iterative revision, and consistent non-sycophantic responses—each demanding the simultaneous exercise of instruction following, context tracking, and reasoning.
@@ 2.2.1 Conversational Engagements in Roleplay
Early persona-grounded dialogue systems aimed at maintaining consistent character personas over multi-turn conversations using architectures like memory networks and transformers. Significant contributions include the persona embeddings introduced by Li et al. (2016), the personalized memory networks developed by Kottur et al. (2017), and notably, the PersonaChat dataset and memory-based models from Zhang et al. (2018), which established foundational benchmarks for persona consistency. These initial works primarily trained models from scratch, facing challenges in sustained persona adherence across interactions. The recent survey by Chen et al. (2024g) extensively covers role-playing before and after the advent of LLMs. Within the scope of this survey, we focus specifically on role-playing interactions under multi-turn LLM settings.
@@
Early LLMs demonstrated an ability to impersonate user-defined personas through prompting alone. Users could supply descriptions like “You are a wise old wizard…” in a system or context prompt, and models like GPT-3 would attempt to respond “in character.” PersonaLLM (Jiang et al., 2024a) introduces a benchmark for personalization and highlighted that simply prefixing instructions with high-level persona descriptions yields limited diversity; instead, it proposed simulating nuanced user preferences via prompt-based reward models. It showed that prompting can move beyond trivial traits to tailor outputs to idiosyncratic user needs. Similarly, CharacterChat (Tu et al., 2023) uses role-playing prompts with behavior presets and dynamic memory to maintain a character’s persona over long conversations. It constructed an MBTI-based persona bank and prompted ChatGPT to produce dialogues between a “seeker” and a compatible “supporter,” injecting preset behavioral tendencies and retrieving context-specific memory each turn to keep interactions coherent.
Beyond persona style, prompting has been used to improve reasoning. Role-play prompting (Kong et al., 2024a) shows that instructing an LLM to “pretend to be” a domain expert can implicitly trigger step-by-step reasoning. In zero-shot settings across 12 tasks, prompting models with a role (e.g. “You are an excellent math teacher…”) led to significantly higher accuracy than a vanilla prompt. The method uses a two-stage prompting framework: first having the model generate an “immersive” backstory or persona acknowledgement, then using that along with the query. These studies collectively demonstrate prompting-based role-play as a powerful lever: it can induce consistent persona adherence (persona and behavior presets) and even boost cognitive performance (reasoning via expert roles) without any parameter updates.
@@
To achieve more robust in-character behavior, researchers introduced instruction tuning and fine-tuning with role-playing data. PIPPA (Gosling et al., 2023) releases a partially synthetic corpus of over 1 million role-play messages, crowdsourced from an online community of role-play enthusiasts. By fine-tuning on PIPPA’s diverse persona-conditioned conversations, small LLMs dramatically improved at staying in character, underscoring that sheer volume and diversity of persona-rich dialogues can teach consistent role-playing. UltraChat (Ding et al., 2023) constructs 1.5M multi-turn dialogues via self-chat with GPT-4, covering broad topics and user types. Fine-tuning LLaMA on this yielded UltraLLaMA, which surpassed previous open models in general conversation quality (including user engagement and coherence). Other data efforts target specific role-play domains: PRODIGy (Occhipinti et al., 2024) builds a dialogue dataset from movie scripts aligned with detailed character profiles (biographies, personality traits). Fine-tuning models on these profile-grounded movie dialogues significantly improved consistency when emulating those characters – e.g. including a character’s backstory and speaking style led to higher human preference for in-character responses.
@@
A parallel direction creates models specialized for particular characters or customizable personas. ChatHaruhi (Li et al., 2023b) focuses on anime characters, compiling 54k dialogues for 32 characters by combining original script lines with simulated conversations. The model fine-tuned on this data, augmented with a “memory” of past events for each character, was able to “revive” characters like Haruhi Suzumiya, accurately quoting lore and personality in new interactions.In another example, CharacterGLM (Zhou et al., 2024a) builds on the Chinese GLM model to allow explicit profile injection for any character: it fine-tuned ChatGLM variants on a corpus of dialogues with richly annotated character profiles (covering identity, style, relationships) and achieved state-of-the-art human-likeness and consistency for customized personas. RoleCraft-GLM (Tao et al., 2024) further extends this concept by crafting original non-celebrity personas with emotional depth and fine-tuning a model on dialogues involving those characters. This yielded more nuanced emotional consistency, validating that meticulous character development during fine-tuning yields agents that are engaging and lifelike in their persona.
@@
Recent work also explores fine-tuning strategies for role consistency. Ditto (Lu et al., 2024) introduces a self-alignment pipeline in which the model generates its own role-play dialogues for 4,000 distinct characters and trains on them. By leveraging the model’s internal knowledge to produce training data—with feedback to ensure character distinctiveness—Ditto achieves strong persona fidelity across a wide range of roles, outperforming fine-tuned baselines on a role-play benchmark. Notably, Ditto frames role-play generation as a reading-comprehension task to avoid style collapse, distinguishing it from UltraChat’s broader synthetic data approach. CharacterLLM (Shao et al., 2023) follows a complementary pipeline: for each target character—particularly historical figures—Wikipedia biographies and documents are used to prompt an LLM to produce character-specific dialogues, and fine-tuning on this synthetic data enables a single agent to robustly portray many roles, effectively transferring factual knowledge into conversational skill. Together, these works reflect a broader trend in role-play fine-tuning from large-scale data harvesting toward increasingly sophisticated data generation and specialization strategies.
@@
Even with instruction tuning, a given model has limits in the number of distinct characters or styles it can perfectly emulate. Thus, a key theme is personalization: adapting an LLM to a new persona with minimal data or effort. Recent research has explored parameter-efficient tuning modules that allow rapid persona swapping without retraining the entire model. PersonaPKT (Han et al., 2023) proposes representing each persona as a continuous embedding vector that can be learned from a small set of that user’s dialogues. By keeping the pre-trained model fixed and only training a tiny persona-specific vector (less than 0.1% of parameters), PersonaPKT efficiently imbues the model with that user’s speaking style and preferences. In a related vein, PPlug (Liu et al., 2025b) introduced a plug-and-play user encoder that on-the-fly computes a user embedding from their conversation history . Instead of a static learned vector, it employs a small model to read a user’s past messages and output a “personal embedding” summarizing their quirks and facts. This embedding is then prepended to the LLM’s input to personalize the response. Such design lets the LLM be dynamically personalized each turn based on context, and experiments showed significant gains in personalization metrics across tasks.
As LLMs are used to power multiple characters simultaneously, it becomes important to switch personas quickly and even maintain many personas at once. Neeko (Yu et al., 2024b) addresses this with a dynamic LoRA (Low-Rank Adapter) framework for multi-character role-play. It pre-trains a separate LoRA module for each character and uses a gating network to activate the appropriate one based on the dialogue context. This incremental ability means an unlimited number of personas can gradually accrue, each encapsulated in a plug-in adapter. An agent can seamlessly swap roles by toggling adapters, which showed superior consistency when one model needed to portray many characters in a group chat. Another challenge is preserving persona over multiple dialogue rounds. Standard fine-tuning often splits dialogues into independent turns, which can break character memory. MIDI-Tuning (Wang et al., 2024c) proposes to explicitly model the user and system roles with separate adapters and a round-level state. In their framework, an LLM-based agent is trained by alternating between a “user adapter” (processing user utterances) and an “agent adapter” (generating responses), carrying hidden state forward through turns.
@@
While supervised learning can teach a model a persona, it doesn’t explicitly punish lapses. Reinforcement learning (RL) provides a way to directly optimize consistency and other long-horizon behaviors. Shea and Yu (2023) demonstrates this by applying offline RL to a dialogue model for persona consistency. They took an existing high-quality chatbot and defined a reward that penalizes persona breaks (e.g. contradicting provided profile or previous statements) and rewards in-character responses. Rather than interact with humans live, they performed RL on a static dataset of conversations – adjusting the model’s policy to maximize the persona consistency reward while leveraging off-policy data. The result was a chatbot that, in human evaluation, more reliably adhered to its given persona description and avoided contradictions compared to its purely supervised counterpart. This work bridged the gap between static fine-tuning and RLHF: it shows one can inexpensively refine a model to be more in-character by offline RL on existing dialogues, getting some benefits of RL (direct control of behavior) without an expensive online loop.
Another aspect of multi-turn role-play that benefits from RL-like thinking is maintaining long-term coherence. COMEDY (Chen et al., 2025c) approaches this via a compressive memory mechanism that can be seen as the model “reinforcing” important memory content over a conversation. Instead of a traditional retrieval pipeline, COMEDY has the LLM periodically summarize and compress the dialogue history (including user persona hints and past events) into a concise memo, which is fed back into itself for future responses. This one-model architecture learns through supervised fine-tuning to generate useful summaries (e.g. remembering the user’s preferences or the agent’s own backstory) and to consult them when answering. While not an RL in algorithm, COMEDY’s design implicitly optimizes a long-term reward: the compressed memory serves to avoid contradictions and boring repetition, much like an RL agent maximizing a reward for user engagement would learn to recall relevant facts.
@@
One line of work focuses on general personalization. The LaMP benchmark (Salemi et al., 2024) evaluates whether LLMs can adapt to user-specific profiles across a range of tasks—such as rewriting text in a personalized tone or classifying based on individual preferences. While not limited to dialogue, LaMP highlights the broader need for systems that understand and leverage identity cues, and confirms that retrieval-based methods are especially effective for on-the-fly personalization.
@@
A more targeted category of benchmarks examines character-specific fidelity. CharacterEval (Tu et al., 2024) is a Chinese benchmark featuring 77 characters from novels, each with multi-turn dialogues and detailed profiles. It defines 13 metrics across four dimensions – including character consistency, behavior realism, and conversational quality – and provides human and model-based evaluation for each. CharacterEval revealed that even GPT-4, when role-playing in Chinese, could be outperformed by fine-tuned local models on consistency, indicating that specialized training made a measurable difference. On the English side, RoleEval (Shen et al., 2023a) poses factual and commonsense questions about 300 well-known characters, assessing whether models accurately retain and apply character-specific knowledge. Results show that global models like GPT-4 excel with internationally known figures, while locally fine-tuned models do better with culturally specific roles. Meanwhile, TimeChara (Ahn et al., 2024) explores a new dimension—temporal consistency—by checking if a model role-playing a character at a given point in a story inadvertently leaks future events. Even the strongest models often violate timeline boundaries, suggesting the need for narrative-aware mechanisms to constrain temporal knowledge. SimulBench (Jia et al., 2025) is a benchmark assessing LLM performance in interactive simulation scenarios – imaginative, role-playing and tool-use tasks that unfold over multiple turns. The benchmark includes tasks like acting as a Linux terminal, playing text-based games, and complex, long-horizon simulations requiring dynamic interaction with a user.
@@
Moving from factual to psychological evaluation, InCharacter (Wang et al., 2024g) assesses if a role-playing agent truly internalizes a character’s personality. It uses an interview-style personality test: the agent (in role) is asked a battery of questions akin to a psychological survey, and its answers are compared to the expected personality profile of the character. Complementing this, RoleInteract (Chen et al., 2024d) evaluates social interaction skills at two levels: one-on-one conversation quality (empathy, politeness, etc.) and group dynamics (how well an agent plays its role in a multi-agent conversation). RoleInteract includes 500 characters and diverse scenarios (e.g., an office meeting with several personas). One finding was that some agents that excel in bilateral chat struggled in group settings – sometimes a normally consistent character would conform or get sidetracked when other AI characters were present, indicating influence and social pressure effects.
@@
A different angle on evaluation is to test how well models understand a character from source material. Yuan et al. (2024) argue that a truly aligned role-play model should be able to read a narrative and produce a coherent character profile. They created the CROSS dataset of expert-written character profiles (covering attributes, relationships, events, and personality) for characters in novels. Models are evaluated on how well they can generate similar profiles after “reading” the novel (or being given chapters as input). In addition to intrinsic metrics (overlap with the expert profile), they evaluate extrinsically via a motivation recognition task: given a scenario from the story, does the model’s profile help it answer why the character acted a certain way. Such evaluation drives home that role-play is not only about output style, but also about the model’s internal model of the character.